### v1

@@ Abstract
Colonoscopic Polyp Re-Identification aims to match the same polyp from a large gallery with images from different views taken using different cameras and plays an important role in the prevention and treatment of colorectal cancer in computer-aided diagnosis. However, traditional methods for object ReID directly adopting CNN models trained on the ImageNet dataset usually produce unsatisfactory retrieval performance on colonoscopic datasets due to the large domain gap. Worsely, these solutions typically learn unimodal modal representations on the basis of visual samples, which fails to explore complementary information from different modalities. To address this challenge, we propose a novel Deep Multimodal Collaborative Learning framework named DMCL for polyp re-identification, which can effectively encourage modality collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Experiments on the standard benchmarks show the benefits of the multimodal setting over state-of-the-art unimodal ReID models, especially when combined with the specialized multimodal fusion strategy.
@@ 2.1 Hand-crafted based Approaches
Traditional research works (Prosser et al, 2010; Zhao et al, 2013, 2014) related to hand-crafted systems for image retrieval task aim to design or learn discriminative representation or pedestrian features. For example, Prosser et al (2010) propose a reformulation of the person re-identification problem as a learning to rank problem. Zhao et al (2013) exploit the pairwise salience distribution relationship between pedestrian images, and solve the person re-identification problem by proposing a salience matching strategy. Besides directly using midlevel color and texture features, some methods (Zhao et al, 2014; Xiang et al, 2020a) also explore different discriminative abilities of local patches for better discriminative power and generalization ability. Unfortunately, these handcrafted feature based approaches always fail to produce competitive results on large-scale datasets. The main reason is that these early works are mostly based on heuristic design, and thus they could not learn optimal discriminative features on current large-scale datasets.
@@ 3.2 Our Proposed DMCL network
Image Encoder. Image encoder is adopted as the visual feature extractor in our DMCL framework. To be more specific, we adopt the ResNet-50 (He et al, 2016) as the backbone for the image encoder module. Benefit from the merits of it’s 50-layer deep network and residual connection design, ResNet-50 network can effectively address the problem of gradient vanishing, enabling the model to deeply learn image features. On the other hand, ResNet-50 can extract rich details and global structures from images by stacking multiple residual blocks to form efficient feature encoding, which exhibits powerful performance and generalization ability in image encoding tasks. To be more specific, given a 224 $\times$ 224 $\times$ 3-sized polyp image, the main goal of the image encoder is to encode the input into a $I^{1\times 2048}$-sized vector for subsequent feature fusion and processing, which provides a strong foundation for downstream polyp ReID task.
@@
Text Encoder. As for the text encoder module, we employ the BERT model (Devlin et al, 2018) to encode the corresponding text description of a polyp. In essence, BERT model is a pre-trained language model based on the Transformer architecture, which learns language representations through masked language modeling and next sentence prediction tasks on a large amount of unlabeled text. In our work, given a text description of the polyp with 5 characters, BERT will encode it into a 5 $\times$ 768-dimensions vector matrix for subsequent feature fusion and processing. The bidirectional encoder design of BERT enables it to capture contextual information of words, generating more accurate text representations.
@@
Dynamic Multimodal Feature Fusion Strategy. The multimodal feature fusion component is a crucial part of our proposed deep collaborative training framework. In the general scenarios, there exists the obvious data discrepancy between image and text in our dataset, which inevitably has a negative impacts on the model’s performance. To overcome this issue, we creatively introduce a multimodal fusion strategy based on the self-attention mechanism. To be more specific, we obtain the image encoding of $I$ after passing through the image encoder, which is then dimensionally reduced from the original 2048-dimensions to 768-dimensions. Similarly, we can also obtain the text encoding through the text encoder, denoted as
$T=[T_{1},T_{2},T_{3},T_{4},T_{5}]$ (1)
@@
where $T$ denotes the text encoding, and $T_{n}$ represents the number index of character in the polyp’s text description. Subsequently, we concatenate the image encoding $I$ and text encoding $T$ to formulate a new encoding of fusion feature, which can be represented as:
$K=[I,T_{1},T_{2},T_{3},T_{4},T_{5}]$ (2)
@@
where $K$ represents the multimodal features after fusion of samples, which is generated by the multimodal fusion component. Remarkably, multimodal fusion component consists of multiple self-attention layers to accomplish the fusion of multimodal features. After passing through three linear layers, these three vectors are then inputted into the self-attention module. The output from the self-attention module is:
@@
where $Q$, $K$, $V$ represents the Query, Key, and Value respectively, and $d$ denotes the dimension of the input vector. After modal fusion, we can obtain the new updated features:
$K_{\text{f}}=\left[I_{f0},T_{f1},T_{f2},T_{f3},T_{f4},T_{f5}\right]$ (4)
Generally speaking, visual feature is of paramount importance for the task of polyp re-identification. Inspired by this, we choose the feature of $I_{f0}$ at the corresponding position of the original image feature, which can be adopted as the testing feature of the corresponding polyp in the testing phase for downstream polyp ReID task.
Algorithm 1 Procedure of our collaborative multimodal training system.
@@
2:Visual sample $x_{v}$ from training dataset $\mathcal{D}$;
3:Textual sample $x_{t}$ from trainig dataset $\mathcal{D}$;
4:Maximum training iterations $n$;
5:Optimized polyp ReID model $\theta_{opt}$;
6:for $iter\leq n$ do
7:     Randomly select a sample $x_{v}$ from training dataset set $\mathcal{D}$;
8:     Randomly select a sample $x_{t}$ from training dataset set $\mathcal{D}$;
9:     Performing visual feature extraction with $x_{v}$;
10:     Performing text feature extraction with $x_{t}$;
11:     Concatenate the image encoding $I$ and text encoding $T$;
12:end for
13:Optimizing model $\theta_{opt}$ by the feature of $K_{\text{f}}$;
14:Performing testing with visual feature and textual feature with model $\theta_{opt}$;
@@ 3.3 Dynamic Network Updating
In essence, many previous works (Xiang et al, 2023a; Luo et al, 2019; Xiang et al, 2020b) have been found that performing training with multiple losses has great potential to learn a robust and generalizable ReID model. During the training process, we adopt a identification loss (Zheng et al, 2017a) and triplet loss function (Hermans et al, 2017) as the optimization metric for our DMCL model. To be more specific, the triplet-loss function aims to reduce the feature distance between similar polyp images while increasing the feature distance between different polyp images. The classification-loss function, on the other hand, is to enhance the model’s ability to recognize and classify polyp categories.
@@
$\mathcal{L}_{ID}=-\frac{1}{\text{ $M_{batch}$ }}\sum_{j=1}^{\text{ $M_{batch}$}}\sum_{i=1}^{N}y_{ji}\log\hat{y_{ji}}$ (5)
where $M_{batch}$ is the number of labeled training images in a batch, $\hat{y_{ji}}$ is the predicted probability of the input belonging to ground-truth class $y_{ji}$. For a triplet couple {$x_{a}$, $x_{p}$, $x_{n}$}, the soft-marginal triplet loss $\mathcal{L}_{triplet}$ can be calculated as:
@@
where $x_{a}$, $x_{p}$ and $x_{n}$ denote the anchor image, positive sample and negative sample respectively. As for the triplet selection, we randomly select different polyps from $16$ patients and sample $4$ images or texts from image or text modality for each patient to form the triplet couple, which can greatly help to construct the embedding space for multimodal representation learning.
@@
To this end, this multimodal testing scheme fully considers the information from both images and texts, providing stronger support for clinical colonoscopy polyp examinations and potentially significantly improving model’s performance.
@@ 4.1 Datasets and Evaluation Metric
4.2 Implementation details
In our experiment, ResNet-50 (He et al, 2016) is regarded as the backbone with no bells and whistles during visual feature extraction. Following the training procedure in (Tong et al, 2022), we adopt the common methods such as random flipping and random cropping for data augmentation and employ the Adam optimizer with a weight decay co-efficient of $1\times 10^{-5}$ and $1\times 10^{-7}$ for parameter optimization. During the training, we adopt the ID loss and triplet loss functions to train the model for 180 iterations, and the cosine distance is also adopted to calculate the similarity of polyp features in the dataset for the task of polyp re-identification. Please note that the text information is also employed during testing phase. In addition, the batch size $M_{batch}$ for training is set to 64. All the experiments are performed on PyTorch framework with one Nvidia GeForce RTX 2080Ti GPU on a server equipped with an Intel Xeon Gold 6130T CPU.
4.3 Comparison with State-of-the-Arts
[…]
@@ 4.3 Comparison with State-of-the-Arts
Person Re-Identification. To further prove the effectiveness of our method on other related object ReID tasks, we also compare our DMCL with existing methods in Table 2. we can easily observe that our method can achieve the state-of-the-art performance on Market-1501, DukeMTMC-reID and CUHK03 datasets with considerable advantages respectively. For example, our DMCL method can achieve a mAP/Rank-1 performance of 92.1% and 96.3% respectively on Market-1501 dataset, leading +1.0% and +0.6% improvement of mAP and Rank-1 accuracy when compared to the second best method NFormer (Wang et al, 2022) and MGN (Wang et al, 2018). Besides, our DMCL method can also obtain the improvement of +1.2% and +1.4% in terms of mAP and Rank-1 accuracy on CUHK03 dataset when compared to the VT-ReID (Xiang et al, 2024b). These experiments strongly demonstrates the priority of our proposed deep multimodal collaborative learning strategy.
@@ 4.4 Ablation Studies
In this experiments, to verify the effectiveness of our deep collaborate training framework, we perform ablation study with multimodal feature from both the qualitative and quantitative perspectives.
Effectiveness of deep multimodal collaborative learning: Firstly, from the quantitative aspect, we evaluate the effectiveness of deep multimodal collaborative learning DMCL framework. As illustrated in Table 3, when adopting our deep multimodal collaborative learning framework, results show that mAP accuracy improves significantly from 25.9% to 46.4% on Colo-Pair dataset with visual-text induction. Additionally, similar improvements can also be easily observed in terms of Rank-1, Rank-5 evaluation metrics, leading to 24.3% and 9.1% respectively. These results prove that multimodal collaborative learning paradigm has a direct impact on downstream polyp ReID task.
@@
Figure 4: Visualization of attention maps with EigenGradCAM: (a) Original images; (b) CNN-based training method without text information; (c) Our deep multimodal collaborative learning method on Colo-Pair dataset. It can be easily observed that semantic pretraining method can capture global context information and more discriminative parts, which are further enhanced in our proposed DMCL method for better performance.
Visualization of Feature Response: To further explain why our deep multimodal collaborative learning strategy works, we perform in-depth analysis of feature response in DMCL method, and also show some qualitative examples of EigenGradCAM (Muhammad and Yeasin, 2020) visualizations in Fig.4. In fact, the Grad-CAM is a package with some methods for Explainable AI for computer vision, which provides attributions to both the inputs and the neurons of intermediate layers, so as to make CNN-based models more transparent by producing visual explanations. Specifically, EigenGradCAM decodes the importance of each feature map to a specific class by analyzing the gradients within the last convolutional layer of a CNN, generating a heatmap that highlights the image regions that contribute most significantly to the prediction result. This heatmap visually represents which parts of the image are most influential in the model’s final decision. Compared with CNN-based training method without text information, we observe that our model attends to relevant image regions or discriminative parts for making decisons, indicating that our deep multimodal collaborate learning framework DMCL can greatly help the model learn more global context information and meaningful visual features with better semantic understanding, which can significantly make our model more robust to perturbations.
@@ 5 Conclusion
This study further investigates the possibility of applying deep multimodal collaborate learning for polyp retrieval task with visual-text dataset, and then propose a simple but effective multimodal representation learning network named DMCL with polyp datasets to improve the performance of polyp Re-ID event. To further enhance the robustness of Re-ID model, a multimodal representation learning strategy is introduced to translate the data of different modalities into the same feature space in the pre-training stage. During the fine-tuning stage, a realistic dataset is adopted to fine-tune the pre-trained Re-ID model for distribution alignment with real-world. Comprehensive experiments on standard benchmark demonstrate that our method can achieve the state-of-the-art performance in polyp Re-ID task. In the future, we will explore the interpretability of this method and apply it to other related computer vision tasks, e.g. polyp detection and segmentation.

### v2

@@ Abstract
Colonoscopic Polyp Re-Identification aims to match the same polyp from a large gallery with images from different views taken using different cameras, which plays an important role in the prevention and treatment of colorectal cancer in computer-aided diagnosis. However, traditional methods for object ReID directly adopting CNN models trained on the ImageNet dataset usually produce unsatisfactory retrieval performance on colonoscopic datasets due to the large domain gap. Worsely, these solutions typically learn unimodal modal representations on the basis of visual samples, which fails to explore complementary information from other different modalities. To address this challenge, we propose a novel Deep Multimodal Collaborative Learning framework named DMCL for polyp re-identification, which can effectively encourage modality collaboration and reinforce generalization capability in medical scenarios. On the basis of it, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Experiments on the standard benchmarks show the benefits of the multimodal setting over state-of-the-art unimodal ReID models, especially when combined with the specialized multimodal fusion strategy, from which we have proved that learning representation with multiple-modality can be competitive to methods based on unimodal representation learning. We also hope that our method will shed light on some related researches to move forward, especially for multimodal collaborative learning. The code is publicly available at https://github.com/JeremyXSC/DMCL.
@@ 2.1 Hand-crafted based Approaches
Traditional research works (Prosser et al, 2010; Zhao et al, 2013, 2014) related to hand-crafted systems for image retrieval task aim to design or learn discriminative representation or pedestrian features. For example, Prosser et al (2010) propose a reformulation of the person re-identification problem as a learning to rank problem. Zhao et al (2013) exploit the pairwise salience distribution relationship between pedestrian images, and solve the person re-identification problem by proposing a salience matching strategy. Besides directly using mid-level color and texture features, some methods (Zhao et al, 2014; Xiang et al, 2020a) also explore different discriminative abilities of local patches for robust generalization. Unfortunately, these handcrafted feature based approaches always fail to produce competitive results on large-scale datasets. The main reason is that these early works are mostly based on heuristic design, and thus they could not learn optimal discriminative features on current large-scale datasets.
@@ 3.2 Our Proposed DMCL network
Image Encoder. Image encoder is adopted as the visual feature extractor in our DMCL framework. To be more specific, we adopt the ResNet-50 (He et al, 2016) as the backbone for the image encoder module. Benefit from the merit of residual connection design, ResNet-50 network can effectively address the problem of gradient vanishing, enabling the model to deeply learn image features. On the other hand, our image encoder can extract rich details and global structures to form efficient feature encoding, which exhibits powerful performance and generalization ability in feature extraction. To be more specific, given a 224 $\times$ 224 $\times$ 3-sized polyp image, the main goal of the image encoder is to encode the input into a $I^{1\times 2048}$-sized vector for subsequent feature fusion and processing, which provides a strong foundation for downstream polyp ReID task.
@@
Text Encoder. As for the text encoder module, we employ the BERT model (Devlin et al, 2018) to encode the text description of corresponding polyp. In essence, BERT model is a pre-trained language model based on the Transformer architecture, which learns language representations through masked language modeling and next sentence prediction tasks on a large amount of unlabeled text. In this work, given a text description of the polyp with $n$ characters, BERT will encode it into a $n$$\times$ 768-dimensions vector matrix for subsequent feature fusion and processing, formulated as $[T_{1},T_{2},T_{3},\cdots,T_{n}]^{1\times 768}$. The bidirectional encoder design of BERT enables it to capture contextual information of words, generating more accurate text representations.
@@
Dynamic Multimodal Feature Fusion. In the general scenarios, there exists the obvious data discrepancy between image and text in our dataset, which inevitably has a negative impacts on the model’s performance. To overcome this issue, we creatively introduce a multimodal fusion strategy based on the self-attention mechanism, which is a crucial part of our proposed deep collaborative training framework. To be more specific, we obtain the image encoding of $I$ after passing through the image encoder, whose size is then dimensionally reduced from the original 2048-dimension to 768-dimension. Similarly, we can also obtain the text encoding through the text encoder, which is denoted as:
$T=[T_{1},T_{2},T_{3},\cdots,T_{n}]$ (1)
@@
where $T$ denotes the text encoding, and $T_{n}$ represents the number index of character in the polyp’s text description. Subsequently, we concatenate the image encoding $I$ and text encoding $T$ to formulate a new encoding of fusion feature, which can be formalized as:
$K=[I,T_{1},T_{2},T_{3},\cdots,T_{n}]$ (2)
@@
where $K$ represents the features generated by the multimodal fusion component. Remarkably, multimodal fusion component consists of multiple self-attention layers to accomplish the fusion of visual feature and text feature. After passing through three linear layers, these three vectors are then inputted into the self-attention module. The output from the self-attention module is:
@@
where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ represents the Query, Key, and Value of self-attention mechanism respectively, and $d$ denotes the dimension of the input vector. After modal fusion, we can obtain the new updated features:
$K_{\text{f}}=\left[I_{f0},T_{f1},T_{f2},T_{f3},\ldots,T_{fn}\right]$ (4)
Generally speaking, visual feature is of paramount importance for the task of polyp re-identification. Inspired by this, we choose the feature of $I_{f0}$ at the corresponding position of the original image feature, which can be adopted as the testing feature of the corresponding polyp in the testing phase for downstream polyp ReID task. The whole procedure of our collaborative multimodal training system is illustrated in Algorithm 1.
Algorithm 1 Procedure of our DMCL training system.
@@
2:Visual sample $x_{v}$ from dataset $\mathcal{D}$;
3:Textual sample $x_{t}$ from dataset $\mathcal{D}$;
4:Initialized deep model $\theta$;
5:Training iterations $n$;
6:Optimized polyp ReID model $\theta_{opt}$;
7:for $iter\leq n$ do
8:     Random select $x_{v}$ from $\mathcal{D}$;
9:     Random select $x_{t}$ from $\mathcal{D}$;
10:     Extracting visual feature $I$ with $x_{v}$ by $\theta$;
11:     Extracting text feature $T$ with $x_{t}$ by $\theta$;
12:     Perform multimodal feature fusion: $K_{\text{f}}\leftarrow I+T$;
[…]
14:end for
15:Performing testing with model $\theta_{opt}$;
@@ 3.3 Dynamic Network Updating
In essence, many previous works (Luo et al, 2019; Xiang et al, 2023a; Luo et al, 2019; Xiang et al, 2020b) have been found that performing training with multiple losses has great potential to learn a robust and generalizable ReID model. During the training process, we adopt a identification loss (Zheng et al, 2017a) and triplet loss (Hermans et al, 2017) as the optimization metric for our DMCL model. Specially, the triplet-loss function aims to reduce the feature distance between similar polyp images while increasing the feature distance between different polyp images. The identification loss function, on the other hand, is to enhance the model’s ability to recognize and classify polyp categories.
@@
$\mathcal{L}_{ID}=-\frac{1}{\text{ $M_{batch}$ }}\sum_{i=1}^{\text{ $M_{batch}$}}\sum_{j=1}^{N}y_{ij}\log\hat{y_{ij}}$ (5)
where $M_{batch}$ is the number of labeled training images in a batch, $\hat{y_{ij}}$ is the predicted probability of the input belonging to ground-truth class $y_{ij}$. For a triplet couple {$x_{a}$, $x_{p}$, $x_{n}$}, the soft-marginal triplet loss $\mathcal{L}_{triplet}$ can be calculated as:
@@
where $x_{a}$, $x_{p}$ and $x_{n}$ denote the anchor image, positive sample and negative sample respectively. As for the triplet selection, we randomly select different polyps from $16$ patients and sample $4$ images or texts from image or text modality for each patient to form the triplet couple, which can greatly help to construct the discriminative embedding for multimodal representation learning.
@@
To this end, our deep multimodal collaborative learning scheme fully considers the information from both images and texts, providing stronger support for clinical colonoscopy polyp examinations and potentially improving model’s performance with a clear margin.
@@ 4.1 Datasets and Evaluation Metric
4.2 Implementation details
In our experiment, ResNet-50 (He et al, 2016) is regarded as the backbone with no bells and whistles during visual feature extraction. Following the training procedure in (Tong et al, 2022), we adopt the common methods such as random flipping and random cropping for data augmentation and employ the Adam optimizer with a weight decay co-efficient of $1\times 10^{-5}$ and $1\times 10^{-7}$ for parameter optimization. Besides, we adopt the ID loss and triplet loss functions to train the model for 180 iterations, and the cosine distance is also adopted to calculate the similarity of polyp features in the dataset for the task of polyp re-identification. Please note that the text information is also employed during testing phase. In addition, the batch size $M_{batch}$ for training is set to 64. All the experiments are performed on PyTorch framework with one Nvidia GeForce RTX 2080Ti GPU on a server equipped with an Intel Xeon Gold 6130T CPU.
4.3 Comparison with State-of-the-Arts
[…]
Person Re-Identification. To further prove the effectiveness of our method on other related object ReID tasks, we also compare our DMCL with existing methods in Table 2. we can easily observe that our method can achieve the state-of-the-art performance on Market-1501, DukeMTMC-reID and CUHK03 datasets with considerable advantages respectively. For example, our DMCL method can achieve a mAP/Rank-1 performance of 92.1% and 96.3% respectively on Market-1501 dataset, leading +1.0% and +0.6% improvement of mAP and Rank-1 accuracy when compared to the second best method NFormer (Wang et al, 2022) and MGN (Wang et al, 2018). Besides, our DMCL method can also obtain the improvement of +1.2% and +1.4% in terms of mAP and Rank-1 accuracy on CUHK03 dataset when compared to the VT-ReID (Xiang et al, 2024b). These experiments strongly demonstrates the priority of our proposed deep multimodal collaborative learning framework.
@@ 4.4 Ablation Studies
In this experiments, to verify the effectiveness of our deep multimodal collaborate learning framework, we perform ablation study with multimodal feature from both the qualitative and quantitative perspectives.
Effectiveness of deep multimodal collaborative learning: Firstly, from the quantitative aspect, we evaluate the effectiveness of deep multimodal collaborative learning DMCL framework. As illustrated in Table 3, when adopting our deep multimodal collaborative learning framework, results show that mAP accuracy improves significantly from 25.9% to 46.4% on Colo-Pair dataset with visual-text induction. Additionally, similar improvements can also be easily observed in terms of Rank-1, Rank-5 evaluation metrics, leading to +24.3% and +9.1% respectively. These results prove that multimodal collaborative learning paradigm has a direct impact on downstream polyp ReID task.
@@
Figure 4: Visualization of attention maps with EigenGradCAM: (a) Original images; (b) CNN-based training method without text information; (c) Our deep multimodal collaborative learning method on Colo-Pair dataset. It can be easily observed that semantic pre-training method can capture global context information and more discriminative parts, which are further enhanced in our proposed DMCL method for better performance.
Visualization of Feature Response: To further explain why our deep multimodal collaborative learning strategy works, we perform in-depth analysis of feature response in DMCL method, and also show some qualitative examples of EigenGradCAM (Muhammad and Yeasin, 2020) visualizations in Fig.4. In fact, the Grad-CAM is a package with some methods for Explainable AI for computer vision, which provides attributions to both the inputs and the neurons of intermediate layers, so as to make CNN-based models more transparent by producing visual explanations. Specifically, EigenGradCAM decodes the importance of each feature map to a specific class by analyzing the gradients within the last convolutional layer of a CNN, generating a heatmap that highlights the image regions that contribute most significantly to the prediction result. This heatmap visually represents which parts of the image are most influential in the model’s final decision. As illustrated in Fig.4, compared with CNN-based training method without text information, we observe that our model attends to relevant image regions or discriminative parts for making decison, indicating that our DMCL model can effectively explore more global context information and meaningful visual features with better semantic understanding, which significantly make our model more robust to perturbations.
@@ 5 Conclusion
This study further investigates the possibility of applying multimodal collaborative learning for polyp retrieval task with visual-text dataset, and then propose a simple but effective multimodal representation learning network named DMCL to improve the performance of polyp Re-ID event. To further enhance the robustness of DMCL model, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Comprehensive experiments on standard benchmarks also demonstrate that our method can achieve the state-of-the-art performance in polyp Re-ID task and other image retriveal task. In the future, we will explore the interpretability of this method and apply it to other related computer vision tasks, e.g. polyp detection and segmentation.